OpenStudy (anonymous):

change 2x+5y=-30 to slope- intercept form. State the slope and y-intercept form.

OpenStudy (anonymous):

y=0.5x-6

OpenStudy (anonymous):

sorry, y=-0.5x-6

OpenStudy (hoblos):

y= -2/5 x -6 slope: 2/5 y-intercept: (0,-6)

OpenStudy (anonymous):

can you show me how you got that ?

OpenStudy (hoblos):

2x+5y=-30 => 5y = -2x -30 divide both sides by 5 => y= -2/5 x - 6

OpenStudy (anonymous):

whats the slope ?

OpenStudy (hoblos):

it is the coefficient of x => -2/5

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

:D

OpenStudy (anonymous):

\[|2x-9|+4\le7 \] then use complete sentence to describe the graph in words.